# An Inversion-based Measure of Memorization for Diffusion Models

### v2

@@ Abstract
The past few years have witnessed substantial advances in image generation powered by diffusion models. However, it was shown that diffusion models are vulnerable to training data memorization, raising concerns regarding copyright infringement and privacy invasion. This study delves into a rigorous analysis of memorization in diffusion models. We introduce an inversion-based measure of memorization, InvMM, which searches for a sensitive latent noise distribution accounting for the replication of an image. For accurate estimation of the memorization score, we propose an adaptive algorithm that balances the normality and sensitivity of the inverted distribution. Comprehensive experiments, conducted on both unconditional and text-guided diffusion models, demonstrate that InvMM is capable of detecting heavily memorized images and elucidating the effect of various factors on memorization. Additionally, we discuss how memorization differs from membership. In practice, InvMM serves as a useful tool for model developers to reliably assess the risk of memorization, thereby contributing to the enhancement of trustworthiness and privacy-preserving capabilities of diffusion models.
@@ 1 Introduction
However, large-scale data sets used to train the prevalent DMs, e.g., the open-source image-caption dataset LAION [45], are widely acknowledged to contain content that will raise concerns regarding copyright infringement, privacy invasion, etc. For example, recent reports revealed that LAION could refer to photographers’ work without authorization [18] and contain private medical photographs [2]. With the uncurated data for training, diffusion models are likely to generate content that infringes the creators’ rights or exposes private information [9, 46]. In this work, we focus on memorization in DMs, one of the most critical issues of training data misuse.
@@
Memorization usually manifests itself at sampling time, when a model generates images closely resembling training images. It can turn into a problem both technically and practically. Technically, memorization indicates failure of generalization to some degree because a probabilistic generative model is supposed to produce novel and diversified images rather than simply replicate those in the training set. Practically, memorization might also raise ethical and legal concerns. First, even the model developers are authorized to train their model with protected images, the image owners will never expect their works to be exposed to arbitrary users due to memorization as this would cause indisciplinable dissemination, violating the contextual integrity [36]. Second, in past years, text-to-image models have been facing lawsuits for generating derivative images that mimic the style of artists. Compared to derivative generations whose legality is still in pending [44], exact replication of copyrighted images is undisputedly intolerable. Finally, a series of works [24, 37] have proposed to use synthetic data in place of real data to prevent sharing of private information. For this goal, potential memorization should also be carefully circumvented.
@@
Recent studies [9, 50, 51, 57] confirm the existence of memorization in DMs in a two-stage way. They first randomly draw samples from the model and then detect risky generations of low diversity [9], search for nearest neighbors in the training set [50, 51] or identify abnormal prediction errors [57]. However, conducting a practical analysis of memorization with these approaches still faces challenges. Transforming the memorization of an image to the presence or absence in samples from the model is not a reliable and efficient solution. While certain images do not appear in a set of samples, it does not demonstrate that they are not highly memorized [54]. For DMs with large latent space, adequate samples should be tested to ensure as complete coverage of the latent space as possible, which usually requires up to millions of testings [57, 9]. Furthermore, things will get much more complicated in conditional DMs, e.g, text-to-image generation. The additional conditions expand the input space of DMs, making the memorization of an image depend on the specific condition. Recent studies use training captions to induce the model to emit corresponding highly memorized training images. However, it still remains unknown that whether the other images not triggered by their coupled captions are also to high extent of memorization.
@@
To address these challenges, we propose an Inversion-based Measure of Memorization (InvMM [^1] The source code can be found at https://github.com/Maryeon/InvMM.) for DMs, an continuous instance-level measure that takes an image as input, outputs a scaler quantifying its level of memorization. As shown in Fig. 1, the inversion process locates a sensitive latent noise distribution and optionally another condition distribution for the target image, such that generating with noise and condition drawn from the inverted distributions will produce samples exactly the same as the target image. InvMM is then defined as the minimum KL divergence to the standard Gaussian distribution and estimated through an adaptive optimization algorithm. Compared to existing studies, InvMM is more reliable and efficient by explicitly searching for all potential sensitive inputs and avoiding to collect any conditions (e.g., prompts). It enables us to discover highly memorized images, claim the safety against memorization for normal images, determine the best training setting, etc. In summary, we make the following contributions in this paper:
  • We propose a novel measure of memorization for DMs, that is more reliable and efficient.
  • For accurate estimation of the measure, we provide an adaptive algorithm that alleviates overestimation and underestimation.
  • We validate the effectiveness of our measure through comprehensive experiments across four datasets, encompassing both unconditional and text-guided DMs.
@@ 2.1 Memorization of generative models
Memorization has raised wide attention on various generative models, including GAN [17], VAE [29], language models [7] and DM [49]. There have been studies on training algorithms [43, 4, 22] and evaluation protocols [19] to improve the generalization ability of generative models that would potentially circumvent memorization. Empirical and theoretical studies also find that dataset complexity and size [15, 50], training time [53], training data duplication [10, 59] and model capacity [5, 10] are influential factors of memorization. Several works [53, 58] focus on the latent space (noise space) to analyze memorization. Nagarajan et al. [53] demonstrate that when trained on sufficient random noise vectors, a generator is guaranteed to memorize training data. Webster et al. [58] measure overfitting of GANs via the reconstruction error at the optimal noise vector. We take a similar perspective on the latent space to analyze memorization while aim to reach an instance-level measure of memorization rather than reveal overall properties of models. In addition, Van der Burg et al. [54] also propose an instance-level measure of memorization as the changed likelihood of one sample when removing it from the training set. However, as discussed by them, the measure is not necessarily comparable between datasets, which is important to understand the true level of memorization.
Specifically for diffusion models, recent research [9, 50, 51] demonstrates that memorization occurs in text-to-image diffusion models. Some heavily memorized images are discovered from numerous generated samples by membership inference [9] or searching for the most similar training images with image matching models [50]. Subsequently, Webster [57] provides more efficient attacks to extract training images from text-to-image models. Wen et al. [61] focus on the detection of abnormal prompts that will trigger generation of training images. In response to the disclosed threats, Vyas et al. [55] propose a provable copyright protection method to prevent replication of sensitive training images. The model is trained to align with a safe model that does not take sensitive data for training. Chen et al. [12] prevent training data replication by a modified sampling procedure that dynamically twists the sampling direction far away from any training points. Our method complements existing studies as a effective way to quantify memorization for any sample (not necessarily training samples).
@@ 2.2 Inversion of diffusion models
Inversion techniques in diffusion models are widely studied for image editing [16, 35, 63], copyright protection [62], generated content detection [56, 11], etc. Through inversion, the object, style and concept contained in the source images can be compressed in latent noises or input token embeddings (for text-to-image generation). Then the inverted latent noises or input token embeddings are utilized to generate same/similar images that preserve the desired content.
@@ 4 Inversion-based measure
Given an image $\bm{x}_{0}$, we aim to measure to what extent $\bm{x}_{0}$ is memorized by a diffusion model (DM). We propose to invert the image into a latent noise distribution, such as noise vectors drawn from it will lead to replication of the image $\bm{x}_{0}$. Our memorization measure is established in this inversion process. In this section, we introduce a sensitivity definition of the inverted noise distribution (Sec. 4.1), elaborate the memorization measure (Sec. 4.2) and give an adaptive algorithm for its calulation (Sec. 4.3). Ultimately, we present an example to measure memorization in conditional DM, specifically text-to-image DM (Sec. 4.4).
@@ 4.1 Sensitivity of latent distribution
The diversity of DM generated content lies in the randomness of latent noise vectors. Starting from different noise vectors, the generated content generally vary. In the context of memorization, this motivates us to search for sensitive noise vectors that will replicate the target image. We then figure out “how many” sensitive noise vectors there are and adopt it as a meaningful measure.
Considering the fact that the common latent noise is sampled from a standard Gaussian distribution, we correspondingly assume the sensitive noise vectors follow a distribution $q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ with learnable parameters $\bm{\varphi}$. The sensitivity function $\mathcal{S}(q_{\bm{\varphi}})$ of a noise distribution can be defined as the probability that generated image $G(\bm{\epsilon}_{0})$ obtained from random noise $\bm{\epsilon}_{0}\sim q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ replicates the target image $\bm{x}_{0}$:
@@
The choice of the distance metric depends on specific applications. In applications that accept derivative generation, a strict metric that only detects eidetic generation can be utilized. While under some circumstances even similar generation should be circumvented, a looser metric is needed. In this work, we consider exact replication that the generated image and target image are exactly the same, following previous work [9]. This consideration to some extent reduces the ambiguity of determining the similarity of two images. Following the investigation of Somepalli et al. [50], we implement the metric as the Euclidean distance between embeddings extracted by the pre-trained copy detection model SSCD [38].
@@ 4.2 Quantify memorization
Based on the definition of sensitivity, we propose an Inversion-based Measure of Memorization (InvMM), as the best achievable normality of $q_{\bm{\varphi}}(\bm{\epsilon}|\bm{x})$ under the condition that $\mathcal{S}(q_{\bm{\varphi}})=1$:
$\min_{\bm{\varphi}}D_{\mathrm{KL}}(q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x})\;\|\;p(\bm{\epsilon}_{0}))$
@@
where $p(\bm{\epsilon}_{0})$ is the prior distribution used in DM, which is often selected to be standard Gaussian. The normality of inverted noise distribution is the KL divergence to the prior.
@@
InvMM requires to achieve the best trade-off between improving normality and meeting the sensitivity condition. Different images to varied extent of memorization may reach different trade-off. Figure 2 illustrates the optimal solution compared to suboptimal choices. Overestimate of memorization produces a distribution with good normality but low sensitivity (Fig. 2(b)). Conservative inversion of the latent distribution perfectly meets the sensitivity condition but underestimates memorization (Fig. 2(c)).
@@ 4.3 Adaptive algorithm
We then describe how to solve the inversion problem and introduce an adaptive algorithm to calculate the proposed measure.
Transforming the inversion problem. It is intractable to directly calculate the sensitivity function. For diffusion models, it can be achieved by minimizing the denoising error $l_{de}$ (see Eq. 7) as a proxy. As can be expected, when $l_{de}$ decreases to a certain level, $\mathcal{S}(q_{\bm{\varphi}})$ will approach 1.
Moreover, denote the KL divergence in Eq. 9 by $l_{kl}$, the optimization problem can be solved by Lagrangian method:
$\min_{\bm{\varphi},\lambda}l_{de}(\bm{x}_{0};\bm{\varphi})+\lambda l_{kl}(\bm{x}_{0};\bm{\varphi}),\lambda>0$ (11)
where $\lambda>0$ is a multiplier. Note that we attach the multiplier to $l_{kl}$ to ensure adequate optimization of the denoising error $l_{de}$.
@@
For the noise distribution $q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$, we assume it to be a multivariate Gaussian $\mathcal{N}(\bm{\mu},\bm{\sigma}^{2})$ with learnable mean $\bm{\mu}$ and diagonal variance $\bm{\sigma}^{2}$, i.e., $\bm{\varphi}=(\bm{\mu},\bm{\sigma}^{2})$. Then InvMM has a closed-form expression $\frac{1}{2}\sum_{i}(\bm{\mu}^{2}_{i}+\bm{\sigma}^{2}_{i}-\log\bm{\sigma}^{2}_{i}-1)$. Although the true distribution of sensitive latent noise might not be fully captured by a Gaussian distribution, we choose this modeling for simplicity and find it effective in experiments.
@@
0:  Noise prediction network $\bm{\epsilon}_{\bm{\theta}}$, target image $\bm{x}_{0}$, optimization iteration $S$, observation cycle $C$, weight increment $\delta$, threshold $\xi$, learning rate $\gamma$
@@
17:     return  $\frac{1}{2}\sum_{i}(\bm{\mu}^{2}_{i}+\bm{\sigma}^{2}_{i}-\log\bm{\sigma}^{2}_{i}-1)$
@@
Adaptive weight tuning. To achieve the best trade-off, the weight $\lambda$ should be adjusted appropriately for different samples. However, it cannot be optimized via gradient decent together with $\bm{\varphi}$. As the gradient w.r.t. $\lambda$ is $l_{kl}$, which is always nonnegative, $\lambda$ will gradually decrease to 0 and $l_{kl}$ will be totally ignored. To solve the problem, we propose an adaptive weight tuning algorithm that allows dynamic change of $\lambda$ during optimization, based on an additive increase/multiplicative decrease scheme similar to congestion control of TCP protocol, as in Algorithm 1, line 6-9.
Specifically, we monitor the denoising error $l_{de}$ every $C$ steps during the optimization process and divide the weight $\lambda$ by half if no significant improvement of denoising error is observed in the cycle. Otherwise, $\lambda$ slowly increases by a small increment $\delta$ each step. In this way, $\lambda$ is adaptively adjusted for different samples to balance the normality and denoising error. When the normality is too strong to further reduce the denoising error, thus to meet the sensitivity condition, it is immediately down-weighted. Otherwise, we gradually strengthen it for better normality. We also perform early stop (line 11-13) when the sensitivity condition is reached to avoid unnecessary normality degradation (underestimate). The sensitivity condition is tested via the Monte Carlo method. Random latent vectors are drawn and used to generate samples. If the distance between any of them and the target image is lower than the threshold $\beta$, then the inversion is successful. There are cases that only partial samples replicate that target image in the end. We still consider such cases as insignificant memorization and flag them as unsuccessful.

### v3

@@ Abstract
The past few years have witnessed substantial advances in image generation powered by diffusion models. However, it was shown that diffusion models are susceptible to training data memorization, raising significant concerns regarding copyright infringement and privacy invasion. This study delves into a rigorous analysis of memorization in diffusion models. We introduce InvMM [^1] https://github.com/Maryeon/InvMM, an inversion-based measure of memorization, which is based on inverting a sensitive latent noise distribution accounting for the replication of an image. For accurate estimation of the measure, we propose an adaptive algorithm that balances the normality and sensitivity of the noise distribution. Comprehensive experiments across four datasets, conducted on both unconditional and text-guided diffusion models, demonstrate that InvMM provides a reliable and complete quantification of memorization. Notably, InvMM is commensurable between samples, reveals the true extent of memorization from an adversarial standpoint and implies how memorization differs from membership. In practice, it serves as an auditing tool for developers to reliably assess the risk of memorization, thereby contributing to the enhancement of trustworthiness and privacy-preserving capabilities of diffusion models.
@@ 1 Introduction
However, large-scale datasets used to train the prevalent DMs, e.g., the open-source image-caption dataset LAION [47], are widely acknowledged to contain content that will raise concerns regarding copyright infringement and privacy invasion. For instance, recent reports revealed that LAION could refer to photographers’ works without authorization [18] and include private medical photographs [2]. With the uncurated data for training, diffusion models are likely to generate content that infringes the creators’ rights or exposes private information [9, 48]. In this work, we focus on memorization in DMs, one of the most critical issues related to training data misuse.
@@
Memorization usually manifests itself at sampling time, when a model generates images closely resembling training images. It can turn into a problem both technically and practically. Technically, heavy memorization indicates failure of generalization to some degree because a probabilistic generative model is supposed to produce novel and diversified images rather than frequently replicate those in the training set. Practically, memorization might also raise ethical and legal concerns. First, even the model developers are authorized to train their model with protected images, the image owners will never expect their works to be exposed to arbitrary users due to memorization as this would cause indisciplinable dissemination, violating the contextual integrity [37]. Second, in past years, text-to-image models have been facing lawsuits for generating derivative images that mimic the style of artists. Compared to derivative generations whose legality is still in pending [46], exact replication of copyrighted images is undisputedly intolerable. Finally, research on privacy protection [25, 38] has proposed to use synthetic data in place of real data to prevent sharing of private information. For this goal, potential memorization must be carefully circumvented.
@@
Recent studies [9, 51, 52, 58] have confirmed the existence of some heavily memorized training data in (text-guided) DMs. They either draw random samples from the model and then detect risky generations of low diversity [9], search for nearest neighbors in the training set [51, 52], or directly identify abnormal prompts [58]. However, conducting a practical analysis of memorization with these approaches still faces challenges. Transforming the memorization of an image to the presence or absence in samples from the model is not a reliable solution. While certain images do not appear in a set of samples, it does not demonstrate that they are not highly memorized [55]. For DMs with large latent space, adequate samples should be tested to ensure as complete coverage of the latent space as possible, which could be up to millions of tests [58, 9]. Furthermore, things will get much more complicated in conditional DMs, e.g, text-to-image generation. The additional conditions expand the input space of DMs, making the memorization of an image also depend on the specific condition. Existing studies exploit training captions to discover highly memorized training images or perform memorization detection [61, 42]. However, attributing the memorization risk of an image to its coupled training caption is incomplete. It still remains unknown that whether the other images not triggered by their coupled captions are also to a large extent of memorization and to be triggered by potential prompts.
@@
Our contributions are threefold. (1) To overcome the challenges, we propose an Inversion-based Measure of Memorization (InvMM) for DMs, a continuous instance-level measure that takes an image as input, outputs a scaler quantifying its level of memorization. As shown in Fig. 1, the inversion process locates a sensitive latent noise distribution and optionally another condition distribution for the target image, such that generating with noise and condition drawn from the inverted distributions will produce samples exactly the same as the target image. InvMM is then defined as the minimum KL divergence of the sensitive noise distribution to the standard Gaussian distribution. Compared to existing methods, InvMM is more reliable and complete in that it proactively searches for all potential sensitive inputs and does not rely on limited test cases from the training set (e.g., captions). (2) To calculate the measure, we design an adaptive optimization algorithm motivated by the TCP congestion control protocol, which is demonstrated to be effective to alleviate overestimate and underestimate of memorization. (3) Through comprehensive experiments across four datasets, encompassing both unconditional and text-guided DMs, we present the superiority of InvMM in analyzing memorization. Compared to alternative memorization measures, detection metrics and membership inference metrics, InvMM is commensurable between samples, reflects the true memorization level from an adversarial perspective and clarifies the fundamental difference between memorization and membership. As an auditing tool, it has great potential in discovering highly memorized images, validating safety against memorization for any image, determining appropriate training settings, etc.
@@ 2.1 Memorization of generative models
Memorization has raised wide attention on various generative models, including GAN [17], VAE [30], language models [7] and DM [50]. There have been studies on training algorithms [45, 4, 22] and evaluation protocols [19] to improve the generalization ability of generative models that would potentially circumvent memorization. Empirical and theoretical studies also find that dataset complexity and size [15, 51], training time [54], training data duplication [10, 60] and model capacity [5, 10] are influential factors of memorization. Several works [54, 59] focus on the latent space (noise space) to analyze memorization. Nagarajan et al. [54] demonstrate that when trained on sufficient random noise vectors, a generator is guaranteed to memorize training data. Webster et al. [59] measure overfitting of GANs via the reconstruction error at the optimal noise vector. We take a similar perspective on the latent space to analyze memorization while aim to reach an instance-level measure of memorization rather than reveal overall properties of models. In addition, Van der Burg et al. [55] also propose an instance-level measure of memorization as the changed likelihood of one sample when removing it from the training set. However, the measure is not necessarily comparable between datasets, which is important to understand the true level of memorization.
Specifically for diffusion models, recent research [9, 51, 52] demonstrates that text-to-image diffusion models suffer from memorization. Some heavily memorized images are discovered from numerous generated samples by membership inference [9] or searching for the most similar training images with image matching models [51]. Webster [58] further provides more efficient attacks to extract training images by identifying abnormal prompts. Subsequent works [61, 42] focus on prompt detection that will trigger training image replication and memorization mitigation [12] at inference time. In response to the disclosed threats, Vyas et al. [56] propose a provable copyright protection method to prevent replication of sensitive training images. Our method improves upon existing methods by being commensurable between different samples, basing on images themselves rather than prompts, and working for any samples (not necessarily training samples). A recent work by Jiang et al. [24] also considers prompt-free memorization. However, their method extends Wen et al.’s findings [61] and is restricted to text-guided diffusion models. Our method is established on the property of latent space thus applies to both unconditional and conditional diffusion models.
@@ 2.2 Inversion of diffusion models
Inversion techniques in diffusion models are widely studied for image editing [16, 36, 63], copyright protection [62], generated content detection [57, 11], etc. Through inversion, the object, style and concept contained in the source images can be compressed in latent noises or input token embeddings (for text-to-image models). Then the inverted latent noises or input token embeddings are utilized to generate similar images that preserve the desired content.
@@ 4 Inversion-based measure
Given an image $\bm{x}_{0}$, we aim to measure to what extent $\bm{x}_{0}$ is memorized by a diffusion model (DM). We propose to invert the image into a latent noise distribution, such that noise vectors drawn from it will lead to replication of the image $\bm{x}_{0}$. Our memorization measure is established in this inversion process. In this section, we introduce a sensitivity definition of the inverted noise distribution (Sec. 4.1), elaborate the memorization measure (Sec. 4.2) and give an adaptive algorithm for its calculation (Sec. 4.3). Ultimately, we present an example to measure memorization in conditional DM, specifically text-to-image DM (Sec. 4.4).
@@ 4.1 Sensitivity of latent distribution
The diversity of DM generated content lies in the randomness of latent noise vectors. Starting from different noise vectors, the generated contents generally vary. In the context of memorization, this motivates us to search for sensitive noise vectors that will replicate the target image. We then figure out “how many” sensitive noise vectors there are and adopt it as a meaningful measure.
Considering the fact that the common latent noise is sampled from a standard Gaussian distribution, we correspondingly assume the sensitive noise vectors follow a distribution $q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ with learnable parameters $\bm{\varphi}$. The sensitivity function $\mathcal{S}(q_{\bm{\varphi}})$ of a noise distribution can be defined as the probability that generated image $G(\bm{\epsilon}_{0})$ obtained from random noise $\bm{\epsilon}_{0}\sim q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ replicates the target image:
@@
The choice of the distance metric depends on specific applications. In applications that accept derivative generation, a strict metric that only detects eidetic generation can be utilized. While under some circumstances even similar generation should be circumvented, a looser metric is needed. In this work, we consider exact replication that the generated image and target image are exactly the same. This consideration to some extent reduces the ambiguity of determining the similarity of two images. Following the investigation of Somepalli et al. [51], we implement the metric as the Euclidean distance between embeddings extracted by the pre-trained copy detection model SSCD [39].
@@ 4.2 Quantify memorization
Based on the definition of sensitivity, we propose an Inversion-based Measure of Memorization (InvMM), as the best achievable average normality of $q_{\bm{\varphi}}(\bm{\epsilon}|\bm{x})$ under the condition that $\mathcal{S}(q_{\bm{\varphi}})=1$:
$\min_{\bm{\varphi}}\frac{1}{N}D_{\mathrm{KL}}(q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x})\;\|\;p(\bm{\epsilon}_{0}))$
@@
where $N$ is the dimensionality of latent noise space and $p(\bm{\epsilon}_{0})$ is the prior distribution used in DM, which is often selected to be standard Gaussian. The normality of inverted noise distribution is the KL divergence to the prior. A lower value of InvMM means greater degree of memorization.
@@
InvMM requires to achieve the best trade-off between improving normality and meeting the sensitivity condition. Different images to varied extent of memorization may reach different equilibrium states. Figure 2 illustrates the optimal solution compared to suboptimal choices. Overestimate of memorization produces a distribution with good normality but low sensitivity (Fig. 2(b)). Conservative inversion of the latent distribution perfectly meets the sensitivity condition while underestimates memorization (Fig. 2(c)).
@@ 4.3 Adaptive algorithm
We then introduce an adaptive algorithm to solve the inversion problem.
Transforming the inversion problem. It is intractable to directly calculate the sensitivity condition. For diffusion models, it can be achieved by minimizing the denoising error $l_{de}$ (see Eq. 7) as a proxy. As can be expected, when $l_{de}$ gradually decreases, $\mathcal{S}(q_{\bm{\varphi}})$ will approach 1.
Moreover, denote the average KL divergence in Eq. 9 by $l_{kl}$, the optimization problem can be solved by Lagrangian method:
$\min_{\bm{\varphi},\lambda}l_{de}(\bm{x}_{0};\bm{\varphi})+\lambda l_{kl}(\bm{x}_{0};\bm{\varphi}),\lambda\geq 0$ (11)
where $\lambda\geq 0$ is a multiplier. We attach the multiplier to $l_{kl}$ to ensure adequate optimization of the denoising error $l_{de}$.
@@
For the noise distribution $q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$, we assume it to be a multivariate Gaussian $\mathcal{N}(\bm{\mu},\bm{\sigma}^{2})$ with learnable mean $\bm{\mu}$ and diagonal variance $\bm{\sigma}^{2}$, i.e., $\bm{\varphi}=(\bm{\mu},\bm{\sigma}^{2})$. Then InvMM has a closed-form expression $\frac{1}{2N}\sum_{i}(\bm{\mu}^{2}_{i}+\bm{\sigma}^{2}_{i}-\log\bm{\sigma}^{2}_{i}-1)$. Although the true distribution of sensitive latent noise might not be fully captured by a Gaussian, we choose this modeling for simplicity and find it effective in experiments.
@@
0: Noise prediction network $\bm{\epsilon}_{\bm{\theta}}$, target image $\bm{x}_{0}$, optimization iteration $T$, observation cycle $C$, weight increment $\delta$, threshold $\xi$, learning rate $\gamma$
@@
17:  return  $\frac{1}{2N}\sum_{i}(\bm{\mu}^{2}_{i}+\bm{\sigma}^{2}_{i}-\log\bm{\sigma}^{2}_{i}-1)$
@@
Adaptive weight tuning. To achieve the best trade-off, the weight $\lambda$ should be adjusted appropriately for different samples. However, it cannot be optimized via gradient decent together with $\bm{\varphi}$. As the gradient w.r.t. $\lambda$ is $l_{kl}$, which is always nonnegative, $\lambda$ will gradually decrease to 0 and $l_{kl}$ will be totally ignored. To solve the problem, we propose an algorithm that allows adaptive tuning of $\lambda$ during inversion, motivated by the additive increase/multiplicative decrease scheme in congestion control of TCP protocol, as in Algorithm 1, line 6-9.
Specifically, we monitor the denoising error $l_{de}$ every $C$ steps during the optimization process and divide the weight $\lambda$ by half if no significant improvement of denoising error is observed in the cycle. Otherwise, $\lambda$ slowly increases by a small increment $\delta$ each step. In this way, $\lambda$ is adaptively adjusted for different samples to balance the normality and denoising error. When the normality is too strong to further reduce the denoising error, thus to meet the sensitivity condition, it is immediately down-weighted. Otherwise, we gradually strengthen it for better normality. We also perform early stop (line 11-13) when the sensitivity condition is reached to avoid unnecessary normality degradation. The sensitivity condition is tested via the Monte Carlo method. Random latent vectors are drawn to generate samples. If the distances between all of them and the target image are lower than the threshold $\beta$, then the inversion is successful. There are cases that only partial samples replicate that target image in the end. We still consider such cases as insignificant memorization and flag them as unsuccessful.